[Hallc_running] [New Logentry] Hall C RC Update (Sept 15)

brads at jlab.org brads at jlab.org
Tue Sep 15 13:55:02 EDT 2020

Logentry Text:
<p> </p><p><strong> Tuesday 15 Sept 2020. </strong></p><h1><span id="General_updates" class="mw-headline">General updates</span></h1><p>(See <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020">d2n Run Plan 2020</a> for broader status/overview.)</p><ul><li>Target work during beam studies yesterday was successful<ul><li>Replaced spectrum analyzer mounting plate. <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848155" rel="nofollow">[1]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848175" rel="nofollow">[2]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848241" rel="nofollow">[3]</a></li><li>Measured target field direction. <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848226" rel="nofollow">[4]</a> These measurements are consistent with Bill/Murchhana's earlier measurements. So target field angle is not the driver for lower tha!
 n expected polarization.</li><li>Rotated target to <strong>Transverse/270</strong> -- Done <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848250" rel="nofollow">[5]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848280" rel="nofollow">[6]</a></li><li>Took some parasitic Carbon-Optics runs during CW portion of beam studies in Kin-A and Kin-X</li></ul></li><li>Beam handed back to us in early OWL, and continued production running in Kin-A, Kin-X at transverse/270<ul><li>Alternating between triggers T1 (1hr), T2 (3hrs); then repeating</li></ul></li><li>Lost SL 2K coldbox at 8am Tuesday morning <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848721" rel="nofollow">[7]</a>; minimum 4 hour down, cryo identified the problem but still working on repair (as of 9:15am)<ul><li>Doing QWP target optimization during this down <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848806" rel="nofollow"!
 >[8]</a></li><li>NMR @ ~42% with calibration constant 5.04 (1:!
 13pm) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848818" rel="nofollow">[9]</a> (Note: calibration constant is estimated).</li><li><strong>Still a few more hours (as of 1:30pm)</strong> <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848822" rel="nofollow">[10]</a></li></ul></li><li>Will continue production running in Kin-A, Kin-X at transverse/270 (w/ T1 and T2 triggers) after standard beam profile checks.</li></ul><h2><span id="Notes_for_the_Future" class="mw-headline">Notes for the Future</span></h2><ul><li>Moller measurement planned for Wednesday DAY starting at 8:30am<ul><li>Hall A's Moller will run Wednesday DAY as well. Will need to coordinate on IHWP switches.</li></ul></li><li>Melanie, Murchhana, Chao, Sylvester: please check good electron yields per Coulomb for Kin-A/X runs with T1 vs. T3 (T2)<ul><li>Mostly complete. Early results don't show any surprises..</li></ul></li></ul><ul><li>EPR calibrations cross checks ar!
 e ongoing [Melanie]<ul><li>So far it all looks good.</li><li>Summary log entries: EPR(270, Kin-Z) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3847972" rel="nofollow">[10]</a>, EPR(90, Kin-X) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3848215" rel="nofollow">[11]</a></li></ul></li></ul><ul><li>Bill, Murchhana check compass/field measurements<ul><li>Also checking with Jixie re: modeled fields</li><li>Need to make a write-up that clearly defines all the different coordinate systems used, swapped power supply/polarity issues, etc.</li></ul></li></ul><ul><li><a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020#1-pass_Running">1-pass running</a> transition scheduled for 6am Friday, Sept 18</li></ul><h2><span id="Pending_Issues" class="mw-headline">Pending Issues</span></h2><ul><li>Review and do any final prep/planning for 1-pass running<ul><li>Paddle optimization for 3He Elastic xsec runni!
 ng [Chao?]</li></ul></li></ul><ul><li>Beam energy drifts<ul><li>Should !
 confirm that MCC remembers to shift the energy lock to Hall C when Hall A turns off</li><li>Just something to keep an eye on. Has not been a problem lately.</li><li>Hall C 5-pass polarization goes down as energy (HALLC:p) goes up <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844493" rel="nofollow">[12]</a></li></ul></li></ul><h2><span id="Opportunistic_Jobs.2FAccess" class="mw-headline">Opportunistic Jobs/Access</span></h2><ul><li><dl><dt>Target calibrations / tests (Particularly anything that will be needed at 1-pass!)</dt></dl><ul><li>Just before Moller measurement, and/or other known downtimes</li><li>See <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020#Target_Calibrations_and_Tests">D2n_Run_Plan_2020#Target_Calibrations_and_Tests</a></li></ul></li></ul><h1><span id="Current_Run_Plan" class="mw-headline">Current Run Plan</span></h1><h2><span id="Tuesday_DAY_.E2.86.92_Tuesday_SWING" class="mw-headline">T!
 uesday DAY → Tuesday SWING</span></h2><ul><li>Wait for 2K coldbox repair, and south linac to be restored</li><li>Continue production data in Kin-A, Kin-X with pol 3He target, <strong>Transverse/270</strong>polarization with pNMR/NMR every 5 hrs.<ul><li><strong>NOTE</strong>: Take 1 (1hr) run with <strong>T1</strong>(3/4) trigger; then 3 runs (1hr each) using <strong>T2</strong>(EL_real) trigger; <strong>Then repeat that cycle.</strong><ul><li>Be sure to monitor and set the prescales correctly.</li><li>Put <strong>T1</strong> or <strong>T2</strong> in the CODA comments field for the runs.</li><li>Note the <strong>T1</strong> or <strong>T2</strong> trigger type for each run in the Shift Summary and Run database using the <code>add_run_comment</code> command.</li></ul></li></ul></li></ul><h2><span id="Wednesday_OWL_.E2.86.92_Wednesday_SWING" class="mw-headline">Wednesday OWL → Wednesday SWING</span></h2><ul><li>Moller measurement on Wednesday DAY at 8:30am (Murchanna, Bill!
 )<ul><li>This will also be used to infer polarization at 1-pass (if it !
 fails, we'll try again on Thursday DAY)</li></ul></li><li><strong>Wednesday OWL or DAY during Moller</strong> Move spectrometers to Kin-C, Kin-Z and continue Transverse/270 running there</li><li>Continue production data in Kin-C, Kin-Z with pol 3He target, <strong>Transverse/270</strong>polarization with pNMR/NMR every 5 hrs when Moller complete.<ul><li>All runs for Kin-C and Kin-Z settings use the <strong>T1</strong> trigger <em>only</em></li></ul></li></ul><pre>-- General plan below this point is fixed, but details may be adjusted as beam/target requires --
</pre><h2><span id="Wednesday_OWL_.E2.86.92_Wednesday_SWING_2" class="mw-headline">Wednesday OWL → Wednesday SWING</span></h2><ul><li>Continue production data in Kin-C, Kin-Z with pol 3He target, <strong>Transverse/270</strong>polarization with pNMR/NMR every 5 hrs.<ul><li>All runs for Kin-C and Kin-Z settings use the <strong>T1</strong> trigger <em>only</em></li></ul></li></ul><h2><span id="Thursday_OWL_.E2.86.92_Friday_OWL" class="mw-headline">Thursday OWL → Friday OWL</span></h2><ul><li>Continue production data in Kin-C, Kin-Z with pol 3He target, <strong>Transverse/270</strong>polarization with pNMR/NMR every 5 hrs.<ul><li>All runs for Kin-C and Kin-Z settings use the <strong>T1</strong> trigger <em>only</em></li></ul></li><li><strong>6 am</strong> Move to Controlled Access (w/ RadCon support) to begin transition to 1-pass</li></ul><h2><span id="Friday_DAY" class="mw-headline">Friday DAY</span></h2><ul><li>(Started at 6am) Open target enclosure<ul><li>Install collima!
 tors (JP, RadCon)</li><li>Check on EPR photo-diode status, likely replace it? (JP, RadCon)</li></ul></li><li>Ensure fast raster is reconfigured correctly (Expert: Jones, Gunning)</li><li>Rotate to minimum angles for HMS (11.7) and SHMS (8.5) (S. Lassiter, A. Comer)</li><li>Configure target for Longitudinal/0 or Longitudinal/180 (TBD)</li></ul><ul><li>Close up and wait for MCC to establish (1-pass) beam to the Hall</li></ul><h2><span id="Friday_SWING_.E2.86.92_Monday" class="mw-headline">Friday SWING → Monday</span></h2><ul><li>See task list at <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020#Run_Plan_for_1-pass">Run Plan for 1-pass</a>for more details<ul><li>Initial priority will be PbPt polarization cross check with longitudinal pol. elastics.</li><li>These have lower priority but will likely be pursued while target spins up<ul><li>3He, N2 density measurements</li><li>H2 xsec measurement</li><li>Pressure curves</li></ul></li><!
 /ul></li></ul><p> </p>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3848826
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200915/9bfb68f2/attachment-0001.html>

More information about the Hallc_running mailing list